OpenStudy (genericnoodles):

If D is the midpoint of CE, CD = x+7, and CE = 5x-1, determine x.

OpenStudy (danjs):

If D is the midpoint on that line, shouldnt CD be half the length of the whole thing CE

OpenStudy (danjs):



OpenStudy (genericnoodles):

So I can count x+7 twice basically and make it equal to 5x-1?

OpenStudy (danjs):

CD = 1/2 CE x + 7 = (1/2)*(5x - 1) solve the x value

OpenStudy (danjs):

yes because it says that it is the midpoint, meaning half the length of the whole line segment

OpenStudy (genericnoodles):

Am I able to do 2x+14=5x-1 and solve from there?

OpenStudy (danjs):

yeah same idea, 2*CD = CE 2 halfs equals the whole

OpenStudy (jadeishere):

Great job!!!!!!!! Welcome to Openstudy!!! :D Sorry. not trying to intrude, continue! Bye now :)

OpenStudy (genericnoodles):

Okay great I understand now, that question really tripped me up and I didn't know what to do lol. Thanks a lot man.

OpenStudy (genericnoodles):

It's all good @jadeishere stop by whenever you want!

OpenStudy (danjs):

welcome, draw the picture helps